# Fair Classifiers Without Fair Training: An Influence-Guided Data Sampling Approach

### v1

@@ D.1.3 Compas Dataset
Compas dataset, also known as the Correctional Offender Management Profiling for Alternative Sanctions dataset, is a collection of data related to criminal defendants. It contains information on approximately 6,000 individuals who were assessed for risk of re-offending. The primary task associated with this dataset is predicting whether a defendant will re-offend ($Y=1$) or not ($Y=0$) within a certain time frame after their release. The sensitive attribute is often considered to be race, specifically whether the individual is classified as African American or not. The dataset has been divided into three subsets: training (4,000 instances), validation (1,000 instances), and test (1,000 instances).
@@
In the experiments using Compas dataset, we train a multi-layer neural network with one hidden layer consisting of 64 neurons. The epochs are split into two phases: warm-up epochs (100 epochs) and training epochs (60 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 128. Furthermore, the default values for learning rate, momentum, and weight decay are 0.01, 0.9, and 0.0005, respectively. We resample the datasets to balance the class and group membership [Chawla et al., 2002]. The dataset is initially split into training and test sets at an 80-20 ratio. Then, we further split 20% of the training set for initial training, reserving the remaining 80% for sampling. Additionally, 20% of the test set is selected to create a validation set. We use the entire model to calculate prediction influence and evaluate fairness for the dataset examples.
D.2 Full Version of Experimental Results
@@
$\epsilon=0.05$ CelebA - Smiling DP EOp EOd Base(ERM) (0.789 ± 0.003, 0.127 ± 0.007) (0.789 ± 0.003, 0.067 ± 0.003) (0.789 ± 0.003, 0.094 ± 0.008) Random (0.837 ± 0.017, 0.133 ± 0.001) (0.840 ± 0.020, 0.059 ± 0.005) (0.837 ± 0.022, 0.039 ± 0.008) BALD (0.856 ± 0.024, 0.140 ± 0.012) (0.854 ± 0.025, 0.062 ± 0.005) (0.856 ± 0.024, 0.033 ± 0.005) ISAL (0.867 ± 0.022, 0.138 ± 0.010) (0.867 ± 0.022, 0.057 ± 0.007) (0.867 ± 0.021, 0.031 ± 0.001) JTT-20 (0.698 ± 0.018, 0.077 ± 0.006) (0.679 ± 0.013, 0.045 ± 0.002) (0.689 ± 0.017, 0.079 ± 0.007) FIS (0.848 ± 0.025, 0.084 ± 0.055) (0.845 ± 0.034, 0.032 ± 0.007) (0.800 ± 0.098, 0.033 ± 0.010)
$\epsilon=0.05$ CelebA - Attractive DP EOp EOd Base(ERM) (0.664 ± 0.005, 0.251 ± 0.022) (0.658 ± 0.009, 0.179 ± 0.014) (0.664 ± 0.005, 0.181 ± 0.017) Random (0.693 ± 0.002, 0.372 ± 0.006) (0.693 ± 0.012, 0.250 ± 0.011) (0.693 ± 0.002, 0.252 ± 0.008) BALD (0.712 ± 0.015, 0.431 ± 0.020) (0.712 ± 0.016, 0.297 ± 0.008) (0.712 ± 0.016, 0.307 ± 0.007) ISAL (0.713 ± 0.009, 0.410 ± 0.003) (0.713 ± 0.009, 0.275 ± 0.011) (0.712 ± 0.009, 0.296 ± 0.023) JTT-20 (0.621 ± 0.015, 0.094 ± 0.004) (0.605 ± 0.013, 0.083 ± 0.011) (0.617 ± 0.013, 0.086 ± 0.006) FIS (0.644 ± 0.025, 0.200 ± 0.060) (0.674 ± 0.015, 0.112 ± 0.025) (0.677 ± 0.010, 0.159 ± 0.008)
@@
Table 6: The performance results of (test_accuracy, fairness_violation) on the Adult dataset. The sensitive attributes is age.
$\epsilon=0.05$ Income (age) DP EOp EOd Base(ERM) (0.665 ± 0.045, 0.255 ± 0.041) (0.665 ± 0.045, 0.115 ± 0.036) (0.665 ± 0.045, 0.158 ± 0.030) Random (0.765 ± 0.021, 0.209 ± 0.042) (0.758 ± 0.027, 0.127 ± 0.013) (0.764 ± 0.018, 0.133 ± 0.027) BALD (0.767 ± 0.019, 0.203 ± 0.017) (0.703 ± 0.111, 0.117 ± 0.013) (0.763 ± 0.022, 0.128 ± 0.014) ISAL (0.765 ± 0.020, 0.215 ± 0.011) (0.755 ± 0.028, 0.128 ± 0.013) (0.761 ± 0.024, 0.138 ± 0.009) JTT-20 (0.751 ± 0.013, 0.262 ± 0.020) (0.742 ± 0.018, 0.149 ± 0.021) (0.745 ± 0.014, 0.171 ± 0.012) FIS (0.766 ± 0.013, 0.214 ± 0.009) (0.757 ± 0.034, 0.113 ± 0.017) (0.763 ± 0.011, 0.143 ± 0.023)
@@
$\epsilon=0.05$ recidivism DP EOp EOd Base(ERM) (0.675 ± 0.005, 0.333 ± 0.008) (0.675 ± 0.005, 0.267 ± 0.010) (0.675 ± 0.005, 0.284 ± 0.010) Random (0.689 ± 0.007, 0.305 ± 0.023) (0.686 ± 0.016, 0.253 ± 0.035) (0.688 ± 0.006, 0.256 ± 0.023) BALD (0.688 ± 0.011, 0.313 ± 0.012) (0.686 ± 0.015, 0.256 ± 0.031) (0.688 ± 0.011, 0.263 ± 0.011) ISAL (0.697 ± 0.002, 0.308 ± 0.025) (0.698 ± 0.004, 0.274 ± 0.022) (0.697 ± 0.001, 0.260 ± 0.026) JTT-20 (0.646 ± 0.009, 0.240 ± 0.016) (0.630 ± 0.024, 0.141 ± 0.028) (0.646 ± 0.009, 0.200 ± 0.007) FIS (0.690 ± 0.002, 0.299 ± 0.029) (0.694 ± 0.002, 0.241 ± 0.035) (0.698 ± 0.005, 0.252 ± 0.030)
D.3 The Impact of Label Budgets
Exploring the impact of label budgets.
In our study, we examine how varying label budgets $r$ influence the balance between accuracy and fairness. We present the results of test accuracy and fairness disparity across different label budgets on the CelebA, Compas, and Adult datasets. In these experiments, we use the demographics parity (DP) as our fairness metric. For convenience, we maintain a fixed label budget per round, using rounds of label budget allocation to demonstrate its impact. The designated label budgets per round for the CelebA, Compas, and Adult are 256, 128, and 512, respectively. In the following figures, the $x$-axis is both the number of label budget rounds. The $y$-axis for the left and right sub-figures are test accuracy and DP gap, respectively. As observed in Figures 2-4, compared to the three baselines (BALD, JTT-20 and ISAL), our approach substantially reduces the DP gap without sacrificing test accuracy.
@@ D.3 The Impact of Label Budgets
Figure 5: We validate how accurate the first-order estimation of the influence is in comparison to the real influence. The $x$-axis represents the actual influence per sample, and the $y$-axis represents the estimated influence. We observe that while some of the examples are away from the diagonal line (which indicates the estimation is inaccurate), the estimated influence for most of the data samples are very close to their actual influence values.

### v3

@@ E.1.3 Compas dataset
Compas dataset, also known as the Correctional Offender Management Profiling for Alternative Sanctions dataset, is a collection of data related to criminal defendants. It contains information on approximately 6,172 individuals who were assessed for risk of re-offending. The primary task associated with this dataset is predicting whether a defendant will re-offend ($Y=1$) or not ($Y=0$) within a certain time frame after their release. The sensitive attribute is often considered to be race, specifically whether the individual is classified as African American or not.
@@
In the experiments using the Compas dataset, we train a multi-layer neural network with one hidden layer consisting of 64 neurons. The epochs are split into two phases: warm-up epochs (20 epochs) and training epochs (50 epochs). The default label budget per round, which represents the number of solicited data samples, is set to 128. Furthermore, the default values for learning rate, momentum, and weight decay are 0.01, 0.9, and 0.0005, respectively. We resample the datasets to balance the class and group membership [17]. The dataset is initially split into training and test sets at an 80-20 ratio. Then, we further split 20% of the training set for initial training, reserving the remaining 80% for sampling. Additionally, 20% of the test set is selected to create a validation set. We use the entire model to calculate prediction influence and evaluate fairness for the dataset examples.
E.2 Full version of experimental results
@@
$\epsilon=0.05$ CelebA - Smiling
[…]
Base(ERM) (0.789 ± 0.003, 0.127 ± 0.007) (0.789 ± 0.003, 0.067 ± 0.003) (0.789 ± 0.003, 0.094 ± 0.008)
Random (0.837 ± 0.017, 0.133 ± 0.001) (0.840 ± 0.020, 0.059 ± 0.005) (0.837 ± 0.022, 0.039 ± 0.008)
BALD (0.856 ± 0.024, 0.140 ± 0.012) (0.854 ± 0.025, 0.062 ± 0.005) (0.856 ± 0.024, 0.033 ± 0.005)
ISAL (0.867 ± 0.022, 0.138 ± 0.010) (0.867 ± 0.022, 0.057 ± 0.007) (0.867 ± 0.021, 0.031 ± 0.001)
JTT-20 (0.698 ± 0.018, 0.077 ± 0.006) (0.679 ± 0.013, 0.045 ± 0.002) (0.689 ± 0.017, 0.079 ± 0.007)
FIS (0.848 ± 0.025, 0.084 ± 0.055) (0.845 ± 0.034, 0.032 ± 0.007) (0.800 ± 0.098, 0.033 ± 0.010)
$\epsilon=0.05$ CelebA - Attractive
[…]
Base(ERM) (0.664 ± 0.005, 0.251 ± 0.022) (0.658 ± 0.009, 0.179 ± 0.014) (0.664 ± 0.005, 0.181 ± 0.017)
Random (0.693 ± 0.002, 0.372 ± 0.006) (0.693 ± 0.012, 0.250 ± 0.011) (0.693 ± 0.002, 0.252 ± 0.008)
BALD (0.712 ± 0.015, 0.431 ± 0.020) (0.712 ± 0.016, 0.297 ± 0.008) (0.712 ± 0.016, 0.307 ± 0.007)
ISAL (0.713 ± 0.009, 0.410 ± 0.003) (0.713 ± 0.009, 0.275 ± 0.011) (0.712 ± 0.009, 0.296 ± 0.023)
JTT-20 (0.621 ± 0.015, 0.094 ± 0.004) (0.605 ± 0.013, 0.083 ± 0.011) (0.617 ± 0.013, 0.086 ± 0.006)
FIS (0.644 ± 0.025, 0.200 ± 0.060) (0.674 ± 0.015, 0.112 ± 0.025) (0.677 ± 0.010, 0.159 ± 0.008)
@@
Table 3: The performance results of (test_accuracy, fairness_violation) on the Adult dataset. The sensitive attribute is age.
$\epsilon=0.05$ Income (age)
[…]
Base(ERM) (0.665 ± 0.045, 0.255 ± 0.041) (0.665 ± 0.045, 0.115 ± 0.036) (0.665 ± 0.045, 0.158 ± 0.030)
Random (0.765 ± 0.021, 0.209 ± 0.042) (0.758 ± 0.027, 0.127 ± 0.013) (0.764 ± 0.018, 0.133 ± 0.027)
BALD (0.767 ± 0.019, 0.203 ± 0.017) (0.703 ± 0.111, 0.117 ± 0.013) (0.763 ± 0.022, 0.128 ± 0.014)
ISAL (0.765 ± 0.020, 0.215 ± 0.011) (0.755 ± 0.028, 0.128 ± 0.013) (0.761 ± 0.024, 0.138 ± 0.009)
JTT-20 (0.751 ± 0.013, 0.262 ± 0.020) (0.742 ± 0.018, 0.149 ± 0.021) (0.745 ± 0.014, 0.171 ± 0.012)
FIS (0.766 ± 0.013, 0.214 ± 0.009) (0.757 ± 0.034, 0.113 ± 0.017) (0.763 ± 0.011, 0.143 ± 0.023)
@@
$\epsilon=0.05$ Recidivism
[…]
Base(ERM) (0.675 ± 0.005, 0.333 ± 0.008) (0.675 ± 0.005, 0.267 ± 0.010) (0.675 ± 0.005, 0.284 ± 0.010)
Random (0.689 ± 0.007, 0.305 ± 0.023) (0.686 ± 0.016, 0.253 ± 0.035) (0.688 ± 0.006, 0.256 ± 0.023)
BALD (0.688 ± 0.011, 0.313 ± 0.012) (0.686 ± 0.015, 0.256 ± 0.031) (0.688 ± 0.011, 0.263 ± 0.011)
ISAL (0.697 ± 0.002, 0.308 ± 0.025) (0.698 ± 0.004, 0.274 ± 0.022) (0.697 ± 0.001, 0.260 ± 0.026)
JTT-20 (0.646 ± 0.009, 0.240 ± 0.016) (0.630 ± 0.024, 0.141 ± 0.028) (0.646 ± 0.009, 0.200 ± 0.007)
FIS (0.690 ± 0.002, 0.299 ± 0.029) (0.694 ± 0.002, 0.241 ± 0.035) (0.698 ± 0.005, 0.252 ± 0.030)
E.3 Exploring the impact of label budgets
In our study, we examine how varying label budgets $r$ influence the balance between accuracy and fairness. We present the results of test accuracy and fairness disparity across different label budgets on the CelebA, Compas, and Adult datasets. In these experiments, we use the demographics parity (DP) as our fairness metric. For convenience, we maintain a fixed label budget per round, using rounds of label budget allocation to demonstrate its impact. The designated label budgets per round for the CelebA, Compas, and Adult are 256, 128, and 512, respectively. In the following figures, the $x$-axis is both the number of label budget rounds. The $y$-axis for the left and right sub-figures are test accuracy and DP gap, respectively. As observed in Figures 5-7, compared to the three baselines (BALD, JTT-20, and ISAL), our approach substantially reduces the DP gap without sacrificing test accuracy.
Specifically, on the Adult dataset, both accuracy and fairness violation converge to similar numerical values when the budget is lower than 20, suggesting a potential overfitting of the model to insufficient training examples. With a larger budget, our algorithms outperform other baseline methods, achieving higher accuracy and lower demographic disparity.
@@ E.3 Exploring the impact of label budgets
E.4 The role of validation dataset size
In this subsection, we explore the impact of adjusting the validation set size on our algorithm’s performance. We present the test accuracy and fairness disparity across different validation set sizes on the CelebA, Compas, and Adult datasets. Note that the default validation set size for image and tabular datasets is set to 1% and 4% of the whole dataset size, respectively. That is, the default validation set sizes are 1996 (CelebA), 1800 (Adult), and 247 (Compas) instances, respectively. In particular, given the smaller size of the default validation set, the minimum scale of the validation set size is set to $\nicefrac{{1}}{{5}}\times$ (nearly 400 CelebA images). Tables 5 and 6 present the performance results on the CelebA, UCI Adult, and Compas datasets, respectively.
Table 5: The performance results of (test_accuracy, fairness_violation) on the CelebA dataset when the validation set size is reduced to $\nicefrac{{1}}{{2}}\times$ and $\nicefrac{{1}}{{5}}\times$. Our algorithm retains the test accuracy and fairness violation when we vary the validation set size.
$\epsilon=0.05$ CelebA - Smiling CelebA - Attractive
[…]
1×1$\times$1 × (0.848, 0.084) (0.876, 0.031) (0.864, 0.030) (0.680, 0.285) (0.695, 0.148) (0.692, 0.148)
$\nicefrac{{1}}{{2}}\times$ (0.872, 0.105) (0.891, 0.042) (0.880, 0.028) (0.648, 0.249) (0.688, 0.188) (0.678, 0.147)
$\nicefrac{{1}}{{5}}\times$ (0.872, 0.117) (0.863, 0.057) (0.886, 0.028) (0.604, 0.171) (0.707, 0.209) (0.645, 0.145)
$\epsilon=0.05$ CelebA - Young CelebA - Big_Nose
[…]
1×1$\times$1 × (0.766, 0.139) (0.775, 0.043) (0.769, 0.168) (0.771, 0.156) (0.765, 0.129) (0.758, 0.155)
$\nicefrac{{1}}{{2}}\times$ (0.735, 0.093) (0.762, 0.067) (0.769, 0.055) (0.771, 0.054) (0.761, 0.162) (0.748, 0.096)
$\nicefrac{{1}}{{5}}\times$ (0.743, 0.107) (0.780, 0.097) (0.757, 0.166) (0.772, 0.095) (0.750, 0.300) (0.760, 0.156)
Table 6: We examine the performance results of (test_accuracy, fairness_violation) on the tabular datasets (Left: Adult; Right: Compas) when the validation set size is reduced to $\nicefrac{{1}}{{2}}\times$, $\nicefrac{{1}}{{4}}\times$, and $\nicefrac{{1}}{{20}}\times$. We observe that our algorithm still retains the test accuracy and fairness violation when we vary the validation set size.
$\epsilon=0.05$ Adult - Income(Age) Compas - Recidivism
[…]
1×1$\times$1 × (0.757, 0.198) (0.718, 0.124) (0.750, 0.125) (0.690, 0.313) (0.696, 0.249) (0.702, 0.257)
$\nicefrac{{1}}{{2}}\times$ (0.717, 0.259) (0.634, 0.123) (0.736, 0.143) (0.683, 0.270) (0.680, 0.247) (0.693, 0.244)
$\nicefrac{{1}}{{4}}\times$ (0.749, 0.196) (0.750, 0.121) (0.747, 0.137) (0.677, 0.283) (0.682, 0.276) (0.680, 0.244)
$\nicefrac{{1}}{{20}}\times$ (0.721, 0.205) (0.706, 0.148) (0.706, 0.179) (0.689, 0.289) (0.668, 0.236) (0.683, 0.252)
E.5 Benchmarking model performance with validation set enhancements
Note that we resort to an additional validation set for developing FIS. To demonstrate FIS’s advantage at the same levels of information, we introduce a new baseline called Random+Val. This method involves continuing to train the model with a randomly sampled validation set. Specifically, we start with the Random’s last saved checkpoint and train it further using the validation set. In particular, we would incorporate a fairness regularizer with dynamic weight to train the model using validation data, considering its sensitive attributes to reduce fairness disparity. Due to its small size, we limit training to 10 epochs to avoid overfitting. The performance results of Random, Random+Val, and FIS are given in Table 7.
Table 7: Comparative analysis of (test_accuracy, fairness_violation) in the CelebA, Adult and Compas datasets. The table illustrates that even at the same information level (using the validation set to train), FIS can obtain better performances. Similarly, we highlight all the fairer but without sacrificing accuracy results in boldface compared to Random.
$\epsilon=0.05$ CelebA - Smiling CelebA - Attractive
[…]
Random (0.853, 0.132) (0.863, 0.053) (0.861, 0.031) (0.696, 0.367) (0.708, 0.253) (0.696, 0.243)
Random + Val (0.801, 0,115) (0.872, 0.139) (0.879, 0.153) (0.638, 0.199) (0.699, 0.366) (0.699, 0.355)
FIS (0.877, 0.122) (0.886, 0.040) (0.882, 0.023) (0.680, 0.285) (0.695, 0.148) (0.692, 0.148)
$\epsilon=0.05$ Adult - Income(Age) Compas - Recidivism
[…]
Random (0.745, 0.236) (0.729, 0.136) (0.748, 0.151) (0.696, 0.316) (0.698, 0.258) (0.694, 0.269)
Random + Val (0.762, 0.161) (0.743, 0.266) (0.788, 0.131) (0.604, 0.136) (0.623, 0.147) (0.628, 0.153)
FIS (0.751, 0.205) (0.725, 0.130) (0.750, 0.125) (0.688, 0.316) (0.695, 0.250) (0.693, 0.254)
Appendix F Potential privacy leakage from non-demographic and demographic feature correlations
Note that some non-demographic features may correlate with sensitive attributes (demographic information), which could lead to potential privacy leakage issues [82]. However, it is crucial to clarify that our method specifically limits its query to the data. We only query the true labels for a selected subset of unlabeled examples. While correlations between non-demographic variables and sensitive attributes may exist in the underlying data, our method itself does not introduce any additional privacy leakage beyond what is inherent in the original dataset. On the other hand, our work’s primary focus is not on addressing privacy concerns related to demographic information. Rather, our main objective is to reduce fairness disparities while maintaining a favorable utility trade-off. We achieve this without explicitly incorporating any additional sensitive information into the process and without directly engaging with the complex privacy implications of demographic data usage.
To address this potential privacy concern, we can analyze it via differential privacy. Consider a function $\phi(\cdot)$ that maps non-demographic information (feature $\bm{x}$ and label $y$) to a sensitive attribute $s$. Due to insufficient information, $\phi(\cdot)$ is not deterministic, making it unable to precisely estimate $s$. If a deterministic mapping function existed, it would unavoidably lead to privacy leakage. Therefore, we assume that given the same features and label, the function $\phi(\cdot)$ output might vary. For example, in the Adult dataset, $\bm{x}$ includes age and credit history, $y$ is the income, and $s$ is the gender. In practice, both men and women have a probability of earning more than 50K (i.e., $y=1$) given the same age, credit history, etc (same feature $\bm{x}$). Suppose $P(\phi(\bm{x},y)=s\mid S=s,\bm{x},y)\leq 1-\epsilon_{0}$ and $P(\phi(\bm{x},y)=s\mid S=s^{\prime},\bm{x},y)\geq\epsilon_{1}$ for all $\bm{x},y,s,s^{\prime}$ (where $s\neq s^{\prime}$). Here, $S$ represents the sensitive attribute variable, which is unknown to $\phi(\cdot)$. Then, we have
$\frac{\mathbb{P}(\phi(\bm{x},y)=\tilde{s}|S=s,\bm{x},y)}{\mathbb{P}(\phi(\bm{x},y)=\tilde{s}|S=s^{\prime},\bm{x},y)}\leq\frac{\max\mathbb{P}(\phi(\bm{x},y)=\tilde{s}\mid S=s,\bm{x},y)}{\min\mathbb{P}(\phi(\bm{x},y)=\tilde{s}\mid S=s^{\prime},\bm{x},y)}\leq\frac{1-\epsilon_{0}}{\epsilon_{1}}=e^{\varepsilon}.$
where $\varepsilon=\ln\left(\frac{1-\epsilon_{0}}{\epsilon_{1}}\right)$. In practice, if the mapping function is too strong, i.e. $\ln\left(\frac{1-\epsilon_{0}}{\epsilon_{1}}\right)$ is too large, we can add additional noise to reduce their informativeness and therefore better protect privacy.